OpenStudy (anonymous):

insert =, <, or > to make the statement true -3/14....-2/3

myininaya (myininaya):

>

OpenStudy (anonymous):

-3/14 > -2/3 because -3/14 = -9/42 -2/3 = -28/42

OpenStudy (anonymous):

soo its >